the area of a square is numerically 5 more than the perimeter. Find the length of the side
Answer by mananth(16946)

l^2=4l+5
l^2-4l-5=0
l^2-5l+l-5=0
l(l-5)+1(l-5)=0
(l-5)(l+1)=0
so l= 5 , ignoring the negative value of l=-1
m.ananth@hotmail.ca